# On the Selection Stability of Stability Selection and Its Applications

### v1

@@ 3 Datasets
As synthetic data, we consider a dataset with a sample size of $n=50$ that includes $p=500$ predictor variables. In this scenario, the predictor variables $\boldsymbol{x}^{\intercal}_{i}$ are independently drawn from the Normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma\in\operatorname{mat}(p,p)$ has a diagonal of ones and $\sigma_{jk}=\rho^{|j-k|}$ for $j\neq k$ where $\rho\in\{0.2,0.5,0.8\}$. The response variable is linearly dependent solely on the first two predictor variables, characterized by the coefficient vector $\boldsymbol{\beta}=(1.5,1.1,0,\ldots,0)^{\top}$, and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample from the standard Normal distribution $\mathcal{N}(0,1)$.
@@
For the first real example, we use the well-established ‘riboflavin’ dataset, which focuses on the production of riboflavin (vitamin B2) from various Bacillus subtilis. This dataset, provided by the Dutch State Mines Nutritional Products, is accessible via hdi R package (Dezeure et al., 2015). It comprises a single continuous response variable that represents the logarithm of the riboflavin production rate, alongside $p=4,088$ covariates, which correspond to the logarithm of expression levels for 4,088 bacterial genes. The primary objective of analyzing this dataset is to identify genes that influence riboflavin production, with the ultimate goal of genetically engineering bacteria to enhance riboflavin yield. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridized during a feed-batch fermentation process involving different engineered strains and varying fermentation conditions. Bühlmann et al. (2014) employed stability selection with LASSO and identified three genes—LYSC_at, YOAB_at, and YXLD_at—as having a significant impact on the response variable.
@@
As an additional real-world example, we investigate ‘Affymetrix Rat Genome 230 2.0 Array’ microarray data introduced by Scheetz et al. (2006). This dataset comprises $n=120$ twelve-week-old male rats, with expression levels recorded for nearly 32,000 gene probes for each rat. The primary objective of this analysis is to identify the probes most strongly associated with the expression level of the TRIM32 probe (1389163_at), which has been linked to the development of Bardet-Biedl syndrome (Chiang et al., 2006). This genetically heterogeneous disorder affects multiple organ systems, including the retina.
In accordance with the preprocessing steps outlined by Huang et al. (2008), we excluded gene probes with a maximum expression level below the $25$th percentile and those exhibiting an expression range smaller than $2$. This filtering process yielded a refined set of $p=3,083$ gene probes that demonstrated sufficient expression and variability for further analysis.
@@ 4 Results
In this section, we illustrate the efficacy of the proposed methodology through its application to the synthetic and real datasets introduced in Section 3.
@@
Here, we present the results obtained from the application of stable stability selection to synthetic data. We used the GitHub repository associated with Nogueira et al. (2018) to implement their stability estimator and to obtain confidence intervals [^1] https://github.com/nogueirs/JMLR2018.
Figure 1 illustrates the selection stability of the stability selection on a grid of regularization values, applied to the synthetic datasets introduced in Section 3, using three different values of $\rho$ and choosing the number of sub-samples $B=500$. The horizontal dashed red lines indicate the stability thresholds of $0.4$ and $0.75$ as discussed in Section 2.
@@
As illustrated in Figure 1, the regularization value that minimizes the cross-validation error, $\lambda_{\text{min}}$, falls within the poorly stable region. In addition, the regularization value that maintains the cross-validation error within one standard error of $\lambda_{\text{min}}$, referred to as $\lambda_{\text{1se}}$ (Hastie et al., 2009), lies within the poor to intermediate stability regions. In contrast, $\lambda_{\text{stable}}$, introduced in Equation (3), is specifically designed to fall within the region of excellent stability, where applicable.
Notably, in all three correlation scenarios, the three regularization selection methods successfully identify relevant variables with high selection frequencies, with a minimum selection frequency of $0.994$. However, since $\lambda_{\text{stable}}$ is larger than the two other regularization values, it applies a stronger shrinkage to variables. To further examine this in terms of model accuracy, for each correlation scenario, we generate $n^{\prime}=25$ additional test samples from the corresponding distribution described in Section 3. At each iteration of stability selection, we predict the response variable for the test samples using the estimated model and, ultimately, we aggregate all the mean squared error (MSE) values obtained over the corresponding $\lambda$ by averaging.
@@
As demonstrated in Figure 2, across all three correlation scenarios, the improvement in stability when transitioning from $\lambda_{\text{1se}}$ to $\lambda_{\text{stable}}$ far outweighs the reduction in accuracy. This outcome aligns closely with the findings of Nogueira et al. (2018), who noted that “All these observations show that stability can potentially be increased without loss of predictive power” and “We also observe that pursuing stability may help identifying the relevant set of features”.
@@
To demonstrate the convergence of the stability estimator $\hat{\Phi}$, for each correlation scenario, we used $\lambda_{\text{stable}}$ and calculated the stability of the selection matrix $M(\lambda_{\text{stable}})$ through sequential sub-sampling, using Equation (2). The stability values across iterations are shown in Figure 3. The blue shading around the line represents the $95\%$ confidence interval for $\hat{\Phi}$. As can be observed, the interval narrows with increasing iterations, indicating a reduction in the estimator’s uncertainty. Figure 3 reveals that, across all three scenarios, after approximately $B^{*}\approx 200$ iterations, the stability estimator converges, with no significant change thereafter in its value. By monitoring the stability trajectory throughout the process, valuable insight can be gained in determining the optimal cut-off point, that is, the optimal number of sub-samples in terms of stability of the process.
@@
Figure 4 illustrates the stability of stability selection when applied to the riboflavin dataset. As shown in Figure 4(a), $\lambda_{\text{stable}}$ does not exist in this example, as the stability values do not surpass the $0.75$ threshold. Figure 4(b), generated with $\lambda_{\text{stable-1sd}}$, indicates the convergence slightly above $0.2$ after about $200$ iterations. These results imply that stability selection with LASSO exhibits poor selection stability within this dataset. Consequently, the findings presented by Bühlmann et al. (2014) may require careful reconsideration, particularly since the selection frequencies of the three identified genes are relatively low.
@@
Due to the instability of the outcomes, $\lambda_{\text{stable}}$ does not exist in this problem. After $B$ iterations, three probes have selection frequencies greater than $0.5$: probe 1390539_at ($\hat{\Pi}_{\texttt{1390539\_at}}^{\lambda_{\text{stable-1sd}}}=0.640$), probe 1389457_at ($\hat{\Pi}_{\texttt{1389457\_at}}^{\lambda_{\text{stable-1sd}}}=0.570$), and probe 1376747_at ($\hat{\Pi}_{\texttt{1376747\_at}}^{\lambda_{\text{stable-1sd}}}=0.564$).
@@
The results illustrated in figures 4 and 5 indicate that while stability selection is a valuable approach, more can be revealed by focusing on the stability of stability selection. As the complexity of the data (in terms of interdependencies between variables, number of variables subject to selection, etc.) increases, the stability values demonstrate a significant decline. Therefore, it is essential to evaluate the stability of its selections before drawing any conclusions about the importance of variables based on the outcomes of stability selection. These findings emphasize the need to carefully assess the stability of the entire process when using stability selection. Neglecting this evaluation may result in overconfidence in the results obtained, even when they lack stability. Given that such analyses are often underrepresented in the literature on stability selection, domain experts may benefit from carefully considering previously published findings.
@@
The convergence plots for both synthetic and real examples indicate that, regardless of whether the convergence value is high or low, stability values tend to stabilize after $B^{*}\approx 200$ iterations. This observation could provide a useful rule of thumb for determining the required number of sub-samples in the stability selection framework.
@@
Nogueira et al. (2018) suggested that stability and accuracy can be analyzed using the concept of the Pareto front (Pareto, 1896), which identifies regularization values that are not dominated by any other in terms of both criteria. A regularization value is considered Pareto optimal if no other value on the regularization grid offers both higher stability and higher accuracy. We call this approach stability-accuracy selection, which seeks to balance these two metrics. In this paper, we introduced the stable stability selection, which focuses on identifying the regularization value that achieves a high stability of the procedure with the least possible loss in accuracy.
As an experiment, in our synthetic data analysis we consider as a measure of accuracy (-MSE), that is, the negative of the mean squared error. Given that a Pareto optimal solution is not necessarily unique, we select the Pareto solution that maximizes the sum of accuracy and stability. In Figure 6(a), $\lambda_{\text{Pareto}}$ is located close to $\lambda_{\text{stable}}$, while in Figures 6(b) and 6(c), they coincide exactly. Interestingly, $\lambda_{\text{stable}}$ is also a Pareto solution in Figure 6(a). It is now pertinent to discuss the relationship between $\lambda_{\text{stable}}$ and Pareto optimality.
@@
Since the stability curve is non-decreasing prior to $\lambda_{\text{stable}}$, any regularization value $\lambda<\lambda_{\text{stable}}\in\Lambda$ is at most as stable as $\lambda_{\text{stable}}$. Therefore, these values do not dominate $\lambda_{\text{stable}}$ in terms of stability.
Similarly, since the loss function is non-decreasing after $\lambda_{\text{stable}}$, any regularization value $\lambda>\lambda_{\text{stable}}\in\Lambda$ is at most as accurate as $\lambda_{\text{stable}}$. Hence, these values also do not dominate $\lambda_{\text{stable}}$ in terms of accuracy.
Since $\lambda_{\text{stable}}$ is not dominated by any values less than it in terms of stability nor by any values greater than it in terms of accuracy, we conclude that $\lambda_{\text{stable}}$ constitutes a Pareto optimal solution. ∎
@@
The two assumptions in Corollary 1 are both natural and justifiable. Increasing the regularization value is expected to allow the model to achieve an optimal balance of sparsity by shrinking irrelevant variables, thereby enhancing stability. The regularization value $\lambda_{\text{stable}}$ is intended to signify the point where the model attains high stability with minimal loss of accuracy. Consequently, we anticipate that the stability curve exhibits a non-decreasing trend prior to reaching $\lambda_{\text{stable}}$. However, since excessive regularization can lead to underfitting, we expect that the stability curve will flatten or decrease after reaching one or more peaks beyond $\lambda_{\text{stable}}$.
Regarding the second assumption, it is expected that increasing regularization helps the model eliminate irrelevant variables, improving its ability to predict the response variable. The regularization values $\lambda_{\text{min}}$ and $\lambda_{\text{1se}}$ are designed to capture this behavior, marking the points where regularization is most effective in terms of predictive ability. Beyond these points, we anticipate that the loss curve will flatten or increase. Since $\lambda_{\text{stable}}$ is intended to trade a small amount of accuracy for greater stability, we expect $\lambda_{\text{stable}}$ to occur after $\lambda_{\text{1se}}$, that is, $\lambda_{\text{stable}}>\lambda_{\text{1se}}>\lambda_{\text{min}}$, which implies that the loss curve is expected to be non-decreasing after $\lambda_{\text{stable}}$.
Thus, according to Corollary 1, $\lambda_{\text{stable}}$ represents a stability-accuracy solution for the problem of regularization tuning, ensuring high stability while minimizing loss in accuracy.
5 Conclusion
In this paper, we examined the stability of the stability selection framework, which underpins the reproducibility of the results and, consequently, our confidence in them. We further applied this concept to determine the optimal regularization value in terms of stability, facilitating the calibration of the decision-making threshold and the expected number of falsely selected variables by leveraging upper-bounds established in the literature. In addition, we introduced a novel selection criterion based on the optimal regularization value, which is called stable stability selection, that ensures the selection of variables associated with highly stable results with the least possible loss in terms of predictive ability. Under two justifiable assumptions, we demonstrated that the proposed optimal regularization value is Pareto optimal in the Pareto front of stability and accuracy. Lastly, we discussed the convergence of stability values across sequential sub-sampling to identify the optimal number of sub-samples based on the stability of the process.
@@ 5 Conclusion
The source code used for the paper is accessible through the following GitHub repository: https://github.com/MahdiNouraie/Stable-Stability-Selection. Furthermore, the stabplot package, which facilitates the use of the two plots introduced in this paper, is available through https://github.com/MahdiNouraie/stabplot.

### v2

@@ 3 Datasets
As synthetic data, we consider a dataset with a sample size of $n=50$ that includes $p=500$ predictor variables. In this scenario, the predictor variables $\boldsymbol{x}^{\intercal}_{i}$ are independently drawn from the Normal distribution $\mathcal{N}(0,\Sigma)$, where $\Sigma\in\operatorname{mat}(p,p)$ has a diagonal of ones and $\sigma_{jk}=\rho^{|j-k|}$ for $j\neq k$ where $\rho\in\{0.2,0.5,0.8\}$. The response variable is linearly dependent solely on the first two predictor variables, characterized by the coefficient vector $\boldsymbol{\beta}=(1.5,1.1,0,\ldots 0)^{\top}$, and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample from the standard Normal distribution $\mathcal{N}(0,1)$.
@@
For the first real example, we use the well-established ‘riboflavin’ dataset, which focuses on the production of riboflavin (vitamin B2) from various Bacillus subtilis. This dataset, provided by the Dutch State Mines Nutritional Products, is accessible via hdi R package (Dezeure et al., 2015). It comprises a single continuous response variable that represents the logarithm of the riboflavin production rate, alongside $p=4,088$ covariates, which correspond to the logarithm of expression levels for 4,088 bacterial genes. The primary objective of analyzing this dataset is to identify genes that influence riboflavin production, with the ultimate goal of genetically engineering bacteria to enhance riboflavin yield. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridized during a feed-batch fermentation process involving different engineered strains and varying fermentation conditions. Bühlmann et al. (2014) employed stability selection with Lasso and identified three genes—LYSC_at, YOAB_at, and YXLD_at—as relevant genes in this problem.
@@
As an additional real-world example, we investigate ‘Affymetrix Rat Genome 230 2.0 Array’ microarray data introduced by Scheetz et al. (2006). This dataset comprises $n=120$ twelve-week-old male rats, with expression levels recorded for nearly 32,000 gene probes for each rat. The primary objective of this analysis is to identify the probes most strongly associated with the expression level of the TRIM32 probe (1389163_at), which has been linked to the development of Bardet-Biedl syndrome (Chiang et al., 2006). This genetically heterogeneous disorder affects multiple organ systems, including the retina. In accordance with the pre-processing steps outlined by Huang et al. (2008), we excluded gene probes with a maximum expression level below the $25$th percentile and those exhibiting an expression range smaller than $2$. This filtering process yielded a refined set of $p=3,083$ gene probes that demonstrated sufficient expression and variability for further analysis.
@@ 4 Results
In this section, we illustrate the efficacy of the proposed methodology through its application to the synthetic and real datasets introduced in Section 3. We used the GitHub repository associated with Nogueira et al. (2018) to implement their stability estimator and to obtain confidence intervals [^1] https://github.com/nogueirs/JMLR2018.
@@
Figure 1 illustrates the selection stability of the stability selection results on a grid of regularization values, applied to the synthetic datasets introduced in Section 3, using three different values of $\rho$ and choosing the number of sub-samples $B=500$. The horizontal dashed red lines indicate the stability thresholds of $0.4$ and $0.75$ as discussed in Section 2. As illustrated in Figure 1, the regularization value that minimizes the cross-validation error, $\lambda_{\text{min}}$, falls within the poorly stable region. In addition, the regularization value that maintains the cross-validation error within one standard error of $\lambda_{\text{min}}$, referred to as $\lambda_{\text{1se}}$ (Hastie et al., 2009), lies within the poor to intermediate stability regions. In contrast, $\lambda_{\text{stable}}$, introduced in Equation (3), is specifically designed to fall within the region of excellent stability, where applicable.
@@
Notably, in all three correlation scenarios, the three regularization selection methods successfully identified relevant variables with high selection frequencies, with a minimum selection frequency of $0.994$. However, since $\lambda_{\text{stable}}$ is larger than the two other regularization values, it applies a stronger shrinkage to variables. To further examine this in terms of model accuracy, for each correlation scenario, we generate $n^{\prime}=25$ additional test samples from the corresponding distribution described in Section 3. At each iteration of stability selection, we predict the response variable for the test samples using the estimated model and, ultimately, we aggregate all the mean squared error (MSE) values obtained over the corresponding $\lambda$ by averaging.
@@
As demonstrated in Figure 2, across all three correlation scenarios, the improvement in stability when transitioning from $\lambda_{\text{1se}}$ to $\lambda_{\text{stable}}$ far outweighs the reduction in accuracy. This outcome aligns closely with the findings of Nogueira et al. (2018), who noted that “All these observations show that stability can potentially be increased without loss of predictive power”.
@@
To demonstrate the convergence of the stability estimator $\hat{\Phi}(\cdot)$, for each correlation scenario, we used $\lambda_{\text{stable}}$ and calculated the stability of the selection matrix $M(\lambda_{\text{stable}})$ through sequential sub-sampling, using Equation (2). The stability values across iterations are shown in Figure 3. The blue shading around the line represents the $95\%$ confidence interval for $\hat{\Phi}(\cdot)$. As can be observed, the interval narrows with increasing iterations, indicating a reduction in the estimator’s uncertainty. Figure 3 reveals that, across all three scenarios, after approximately $B^{*}\approx 200$ iterations, the stability estimator converges, with no significant change thereafter in its value. By monitoring the stability trajectory throughout the process, valuable insight can be gained in determining the optimal cut-off point, that is, the optimal number of sub-samples in terms of stability of the results.
@@
Figure 4 illustrates the stability of stability selection when applied to the riboflavin dataset. As shown in Figure 4(a), $\lambda_{\text{stable}}$ does not exist in this example, as the stability values do not surpass the $0.75$ threshold. Figure 4(b), generated with $\lambda_{\text{stable-1sd}}$, indicates the convergence slightly above $0.2$ after about $200$ iterations. These results imply that stability selection with Lasso exhibits poor selection stability on this dataset.
@@
Due to the instability of the results, $\lambda_{\text{stable}}$ does not exist in this problem. After $B$ iterations, three probes have selection frequencies greater than $0.5$: probe 1390539_at ($\hat{\Pi}_{\texttt{1390539\_at}}^{\lambda_{\text{stable-1sd}}}=0.640$), probe 1389457_at ($\hat{\Pi}_{\texttt{1389457\_at}}^{\lambda_{\text{stable-1sd}}}=0.570$), and probe 1376747_at ($\hat{\Pi}_{\texttt{1376747\_at}}^{\lambda_{\text{stable-1sd}}}=0.564$).
@@
The results illustrated in figures 4 and 5 indicate that while stability selection is a valuable approach, more can be revealed by focusing on the stability of stability selection results. As the complexity of the data (in terms of interdependencies between variables, number of variables subject to selection, etc.) increases, the stability values demonstrate a significant decline. Therefore, it is essential to evaluate the stability of its selections before drawing any conclusions about the relevance of variables based on the results of stability selection.
@@
The convergence plots for both synthetic and real examples indicate that, regardless of whether the convergence value is high or low, stability values tend to stabilize after $B^{*}\approx 200$ iterations. This observation provides a useful rule of thumb for selecting the required number of sub-samples in the stability selection framework when using the Lasso.
@@
Nogueira et al. (2018) suggested that stability and accuracy can be analyzed using the concept of the Pareto front (Pareto, 1896), which identifies regularization values that are not dominated by any other in terms of both criteria. A regularization value is considered Pareto optimal if no other value on the regularization grid offers both higher stability and higher accuracy. We call this approach stability-accuracy selection, which seeks to balance these two metrics.
As an experiment, in our synthetic data analysis we consider the negative of the mean squared error (-MSE) as a measure of prediction accuracy. Given that a Pareto optimal solution is not necessarily unique, we select the Pareto solution that maximizes the sum of accuracy and stability. In Figure 6(a), $\lambda_{\text{Pareto}}$ is located close to $\lambda_{\text{stable}}$, while in Figures 6(b) and 6(c), they coincide exactly. Interestingly, $\lambda_{\text{stable}}$ is also a Pareto solution in Figure 6(a). It is now pertinent to discuss the relationship between $\lambda_{\text{stable}}$ and Pareto optimality.
@@
Since the stability curve is non-decreasing prior to $\lambda_{\text{stable}}$, any regularization value $\lambda<\lambda_{\text{stable}}\in\Lambda$ is at most as stable as $\lambda_{\text{stable}}$. Therefore, these values do not dominate $\lambda_{\text{stable}}$ in terms of stability. Similarly, since the loss function is non-decreasing after $\lambda_{\text{stable}}$, any regularization value $\lambda>\lambda_{\text{stable}}\in\Lambda$ is at most as accurate as $\lambda_{\text{stable}}$. Hence, these values also do not dominate $\lambda_{\text{stable}}$ in terms of accuracy. Since $\lambda_{\text{stable}}$ is not dominated by any values less than it in terms of stability nor by any values greater than it in terms of accuracy, we conclude that $\lambda_{\text{stable}}$ constitutes a Pareto optimal solution. ∎
@@
The two assumptions in Corollary 1 are both justifiable. Increasing the regularization value is expected to allow the model to achieve an optimal balance of sparsity by shrinking irrelevant variables, thereby enhancing stability. The regularization value $\lambda_{\text{stable}}$ is intended to signify the point where the model attains high stability with minimal loss of accuracy. Consequently, we anticipate that the stability curve exhibits a non-decreasing trend prior to reaching $\lambda_{\text{stable}}$.
Regarding the second assumption, it is expected that increasing regularization helps the model eliminate irrelevant variables, improving its ability to predict the response variable. The regularization values $\lambda_{\text{min}}$ and $\lambda_{\text{1se}}$ are designed to capture this behavior, marking the points where regularization is most effective in terms of predictive ability. Beyond these points, we anticipate that the loss curve will be non-decreasing. Since $\lambda_{\text{stable}}$ is intended to trade a small amount of accuracy for greater stability, we expect $\lambda_{\text{stable}}$ to occur after $\lambda_{\text{1se}}$, that is, $\lambda_{\text{stable}}>\lambda_{\text{1se}}>\lambda_{\text{min}}$, which implies that the loss curve is expected to be non-decreasing after $\lambda_{\text{stable}}$.
Thus, according to Corollary 1, $\lambda_{\text{stable}}$ represents a stability-accuracy solution for the problem of regularization tuning, ensuring high stability while minimizing loss in prediction accuracy.
@@
The source code used for the paper is accessible through the following GitHub repository: https://github.com/MahdiNouraie/Stable-Stability-Selection. Furthermore, the stabplot R package, which facilitates the use of the two plots introduced in this paper, is available through https://github.com/MahdiNouraie/stabplot.